### v1

# Hybrid Video Anomaly Detection for Anomalous Scenarios in Autonomous Driving

Abstract
In autonomous driving, the most challenging scenarios are the ones that can only be detected within their temporal context. Most video anomaly detection approaches focus either on surveillance or traffic accidents, which are only a subfield of autonomous driving. In this work, we present HF2-VADAD, a variation of the HF2-VAD surveillance video anomaly detection method for autonomous driving. We learn a representation of normality from a vehicle’s ego perspective and evaluate pixel-wise anomaly detections in rare and critical scenarios.
[…]
Keywords: Autonomous Driving, Anomaly Detection, Behavioral Anomalies
1 Introduction
In autonomous driving, the detection of anomalies is crucial for ensuring safety and reliability. Video anomaly detection (VAD) focuses on identifying events in video data that deviate from an expected normality. In autonomous driving, the challenges of detection are complicated by factors such as camera movements, ever-changing backgrounds, and rapid changes in vehicle speed.
Many different types of anomalies exist [1, 2, 3], with many approaches trying to detect them [4]. We can distinguish between five key techniques: Reconstruction, prediction, generative modeling, feature extraction, and confidence evaluation [5]. Reconstruction-based methods focus on learning a representation of normality and labeling deviations from this norm [6, 7]. Predictive approaches detect anomalies when future predictions deviate significantly from observations [8, 9]. However, in autonomous driving, most of the anomaly detection methods that are based on raw sensory data focus on atypical objects rather than atypical behaviors. Temporal anomalies are typically detected in object-based trajectory data, assuming perfect perception [10, 11, 12, 13]. The most complex temporal anomalies are anomalous scenarios, which pose a high risk of collision [1].
Figure 1: Exemplary anomaly detection with HF2-VADAD. In the left frame, the vehicle in front performs a rare sudden braking maneuver in simulation, as highlighted in yellow in the top-left graph. The graph shows the anomaly scores per frame, peaking while the vehicle performs the maneuver. The right frame shows pixel-wise anomaly scores for all instances of the class vehicle.
Until recently, mostly accident-focused benchmarks were available to evaluate temporal anomalies in the context of autonomous driving [14]. Given a rise in datasets with novel dynamic anomalies [15, 16, 17], we are interested in the applicability of video anomaly detection methods for autonomous driving. Among recent state-of-the-art methods, we choose HF2-VAD as the representative model [18] and adapt it to the field of autonomous driving. For this purpose, we train a HF2-VADAD model on simulated data representing normality and evaluate it on the challenging AnoVox dataset [17] with anomalous scenarios in the form of sudden braking maneuvers of the lead vehicle. We introduce a pixel-wise evaluation of HF2-VADAD to also localize anomalies.
2 Related Work
There are three typical camera perspectives for anomaly detection methods in the traffic domain. Birds-Eye-View (BEV) and surveillance perspectives come from external data providers, while an ego or dashcam view can be provided by a vehicle itself [14]. Numerous works have focused on specific scenarios, including the detection of atypical activity in urban surveillance, the analysis of irregularities in pedestrian dynamics, and the monitoring of abnormal behavior patterns in crowds [19, 20, 21, 22, 23, 24, 25, 26, 27] as well as street scenes recorded by city cameras that evaluate vehicle behavior from an outside perspective [28, 29].
Especially the VAD subfields traffic anomaly detection [30, 31] and road accident detection [32] have gained much traction. However, only a few methods in the literature specifically analyze behavioral anomalies from the perspective of a vehicle’s onboard cameras. For our comparison in Table 1, we have also excluded frame-wise [33, 34, 35, 36] or supervised [37, 38, 39, 40] methods as well as such that focus on domain-level anomalies [41, 42, 43, 44].
In general video anomaly detection, hybrid methods that combine both reconstruction and prediction methods have emerged as the current state-of-the-art. These methods aim to utilize the strengths of both approaches by leveraging temporal dynamics and reconstruction capabilities to improve anomaly detection [45, 46, 18, 47, 48, 49]. Among these, HF2-VAD by Liu et al. [18], who built upon the work of Park et al. [50], is commonly used in the literature as a benchmark for the state of the art in video anomaly detection [51] and often serves as the foundation for further advancements [52]. It detects unusual events that deviate from normal patterns for surveillance tasks.
The idea of predicting future frames for the purpose of anomaly detection became popular around 2019, where it was demonstrated both for surveillance tasks [8] and autonomous driving [53, 54]. Liu et al. combined future frame predictions with an optical flow [8], and Bolte et al. focused on semantically distinguishing relevant classes from the rest of an image [53]. Yao et al., on the other hand, focused on predicting only the bounding boxes of relevant classes in order to compensate for the motion of the ego vehicle, which makes the prediction of whole frames more complex [54]. Later, Fang et al. built upon their work but removed the need for an odometry estimation for the ego vehicle. Recently, Yao et al. extended their own work and also provided an updated version of Liu’s approach by adding masks for relevant objects, similar to how Bolte approached the topic [55]. Finally, the new DeepAccident dataset by Wang et al. also proposed an accident detection method based on future trajectories and their distances to each other [15]. As shown in Table 1, most of the methods require many additional sub-methods to function, making them rather complex and only recently began to work for high-resolution input images.
Table 1: Behavioral, pixel-wise anomaly detection methods for vehicle camera data in motion.
Method Year High Resolution Required Sub-Methods Anomaly Detection
ROI (BEV) Mask Bbox Flow Odometry Tracking Motion Pred. Graph
AnoPred [8, 54, 55] 2019 Learned Normality
CC-Detector [53] 2019 Learned Normality
SSC-TAD [56] 2022 Learned Normality
FOL [55, 54] 2023 Learned Normality
AnoPredMask [8, 55] 2023 Learned Normality
AccPred [15] 2023 Metric-based
HF2-VADAD [18] 2024 Learned Normality
3 Methodology
As shown in Section 2 and Table 1, existing video anomaly detection methods for autonomous driving are either overly complex or incapable of dealing with high-resolution input data. Based on HF2-VAD [18], our adaption HF2-VADAD is capable of utilizing high-resolution video streams, learns a representation of normality in the domain of autonomous driving and only requires two sub-methods. Thus, it is the least complex method for video anomaly detection in the field of autonomous driving for high-resolution input.
Figure 2: HF2-VADAD: Adaptation and extension of HF2-VAD for autonomous driving. Optical flows ${y}_{1:t}$ and bounding boxes ${x}_{1:t}$ for relevant objects are generated for each input image. ML-MemAE-SC reconstructs the optical flows $\hat{y}_{1:t}$ with memory-modules $M$ to better reconstruct only normal patterns. The CVAE predicts a future frame $\hat{x}_{t+1}$. On this basis, image-wise and localized pixel-wise anomaly scores are generated. Adapted from [18].
HF2-VAD. Here, we provide a brief overview of the hybrid framework HF2-VAD [18] we build upon and refer to the original publication for further details. The framework combines Multi-Level Memory modules in an Autoencoder with Skip Connections (ML-MemAE-SC) for optical flow reconstructions and a Conditional Variational Autoencoder (CVAE) for future frame predictions. The framework is trained on normal data and uses discrepancies in the reconstructions and predictions to identify anomalies. Experiments were conducted with well-established surveillance datasets.
The ML-MemAE-SC uses an autoencoder structure with multiple memory modules at different levels. This structure improves the model’s ability to learn and retain typical data patterns from the training set and distinguish them from anomalies. The integration of skip connections between the encoder and decoder helps to preserve important data that might otherwise be lost during the encoding process. This part focuses on the accurate reconstruction of normal data. The training of the ML-MemAE-SC includes a composite loss function that summarizes reconstruction and entropy losses. The reconstruction loss is calculated based on the Euclidean distance.
The CVAE plays a central role in predicting future images by using conditioned historical data to simulate probable future scenarios. It determines the latent variable $z_{t+1}$ by considering both the current image $x_{t}$ and the optical flow $y_{t,t+1}$ between successive images. It works on the premise that a latent variable $z$, representing the probabilistic state of the following image, can be derived from the immediate past to subsequently influence the generation of $\hat{x}_{t+1}$. It uses the encoded optical flow and the current image to estimate the posterior distribution $p(z_{t+1}|x_{t},y_{t,t+1})$. A sample from this distribution, which represents the probabilistic state of the next image, is then drawn for image generation. During training, this latent variable $z$ is combined with the conditions derived from the encoder $E$ and processed by the decoder $D$ to create the future image $\hat{x}_{t+1}$.
At test time, anomaly detection is performed frame-wise. The frame-wise anomaly score $S_{f}$ is a fusion of the flow reconstruction error $S_{r}$ and the error in the prediction of future images $S_{p}$, standardized by their respective means $\mu$ and standard deviations $\sigma$ from the training samples:
$S_{f}=w_{r}\cdot\frac{S_{r}-\mu_{r}}{\sigma_{r}}+w_{p}\cdot\frac{S_{p}-\mu_{p}}{\sigma_{p}}$ (1)
HF2-VADAD. First, we adapt HF2-VAD to autonomous driving by generating dense, pixel-wise anomaly scores for the whole frame rather than focusing on the content of unlocalized, detected bounding boxes alone, as can be seen in Figure 1. We learn flow reconstructions and flow-guided frame prediction from the perspective of an ego-vehicle rather than from the perspective of a static surveillance camera. Our training dataset was collected with the CARLA simulation engine [57]. Rather than collecting data with the default autopilot, we employ Roach [58] as the driving agent.
To reflect the complexity of real-world traffic, the training dataset features four different cities, each with its own weather and traffic dynamics. The training data includes 38 scenarios with over 30,000 high-resolution frames with a resolution of 600 x 800 pixels. Crucially, the training dataset contains only normal driving scenes so that the model can internalize a baseline of normal traffic behavior, which is crucial for detecting deviations that indicate anomalies in real-world driving situations. Other normally behaving agents, including lead vehicles, are included in the training data.
Flows are generated by FlowNet2.0  [59] to ensure that both static and motion-based features contribute to the training of the model. Bounding boxes are detected with an off-the-shelf object detection model [60] trained on CARLA data.
We first train the ML-MemAE-SC with the extracted flows from the training dataset, then train the CVAE model with the reconstructed flows from the ML-MemAE-SC and the corresponding frames, and finally fine-tune the whole framework. During fine-tuning, the weights of the two models are adjusted together to improve their coordination and the overall accuracy of anomaly detection.
We train the model specifically on foreground objects, i.e., vehicles, extracted from these images with bounding boxes, and construct spatiotemporal cubes (STC) for each object, where the dimensions of these cubes are standardized to 32x32. The use of bounding boxes for each vehicle in the scene allows the model to focus specifically on regions within the image where temporal anomalies in the form of unexpected maneuvers can occur.
Figure 3: Sequential Anomaly Detection: The first image shows the normal driving mode of the lead vehicle, while the last two images show an anomaly with the pixel-wise ground truth overlayed in red. HF2-VADAD successfully detects the unknown maneuver.
In addition to the existing frame-wise anomaly detection, we have integrated a pixel-wise anomaly detection. For this, the anomaly score for each pixel, denoted by $S_{\text{pixel}}$, is derived from a weighted average of robustly scaled mean squared errors (MSE) for optical flow and image prediction:
$S_{\text{pixel}}=w_{\text{r,p}}\cdot m_{\text{r}}+w_{\text{p,p}}\cdot m_{\text{p}}$ (2)
The variables $m_{\text{r}}$ and $m_{\text{p}}$ denote the mean values of the robustly scaled MSEs for the optical flow and image prediction, respectively. The scores are calculated using the weights $w_{\text{r,p}}$ for optical flow and $w_{\text{p,p}}$ for image prediction. Robust scaling is used to standardize the MSEs by reducing the influence of outliers, making the method more reliable for detecting anomalies under different conditions. The averaging process calculates the mean of these scaled values to summarize the error across all pixels. The resulting pixel-wise scores are then used to localize anomalies within the image. A score of 0 is set for all pixels outside the bounding boxes, as shown in Figure 1. In contrast, the frame-wise approach of the original HF2-VAD uses simple summation and normalization based on the training statistics. Specifically, the sum of MSE losses across all dimensions is calculated for both optical flow and image predictions, and these summed values are then normalized using the mean and standard deviation from the training data. In the presence of detections, they add up all $32\times 32$ anomaly patches and choose the largest value as the frame-wise anomaly score. We also report these frame-wise results as AUROC in Section 4.
In contrast, our approach calculates a pixel-wise anomaly score by normalizing and aggregating mean-squared errors of optical flow and frame predictions, then scaling these scores to bounding boxes within each frame. This pixel-wise approach allows for a more granular analysis, focusing on specific areas within frames rather than a single aggregated frame-level score.
4 Evaluation
For the evaluation, we used the recently released AnoVox benchmark [17] with anomalous scenarios in the form of sudden braking maneuvers performed by lead vehicles, as visible in Figure 3. These are common and dangerous situations in road traffic [61] and are not included in our training data, as Figure 4 shows. In the 13 evaluation scenarios, the ego-vehicle follows a lead vehicle along a longer route under different weather and traffic conditions. At certain intervals, the lead vehicle performs a sudden braking maneuver. We extended the available pixel-wise ground truth, as shown in Figure 3 by frame-wise annotations. The annotation of the anomalies started with the unexpected sudden braking of the leading vehicle and ended with its complete stop.
Figure 4: Distribution of braking behavior in training (blue) and test (red) data. High-intensity sudden braking scenarios only occur in the evaluation dataset.
Table 2: Evaluation and ablation studies of HF2-VADAD and ML-MemAE-SC under different scenarios, with best and second-best results highlighted.
Model Domain Weather Finetuned AUROC $\uparrow$ FPR95 $\downarrow$ IoU $\uparrow$
Bounding Boxes: Detection
HF2-VADAD All All 66.98 2.58 48.10
HF2-VADAD City Sunshine 68.04 2.48 48.09
HF2-VADAD City Rain 69.74 2.68 51.30
HF2-VADAD Highway Sunshine 57.22 3.15 60.71
HF2-VADAD Highway Rain 76.60 1.34 41.09
Bounding Boxes: Ground Truth
ML-MemAE-SC All All 67.49 7.80
HF2-VADAD All All 81.01 3.72
HF2-VADAD All All 79.09 2.83
HF2-VADAD City Sunshine 80.74 3.10
HF2-VADAD City Rain 74.08 3.05
HF2-VADAD Highway Sunshine 74.76 3.19
HF2-VADAD Highway Rain 91.74 99.32
In order to evaluate the performance of HF2-VADAD, we provide both quantitative and qualitative insights. Two metrics are used for evaluation: the already used Area Under the Receiver Operation Characteristic (AUROC) for frame-wise evaluations and the False Positive Rate at 95% True Positive Rate (FPR95) for pixel-wise evaluation. The AUROC metric, widely used in the VAD literature [62, 18, 63, 64]], serves as a measure of VAD accuracy by varying the threshold for anomaly assessment, with higher AUROC values indicating better detection performance. Adding FPR95 [65, 66] provides insights into the false-positive rates if nearly all anomalies are correctly detected, with lower values indicating better detection performance. Since our detected bounding boxes do not fully match the existing ground truth, we only evaluate the FPR95 based on the overlapping area and also provide the IoU to demonstrate the quality of the bounding boxes.
4.1 Quantitative Results
As shown in Table 2, we perform various experiments to better understand the performance of HF2-VADAD under varying conditions. Comparing city and highway scenarios, we observe a strong sensitivity of the model to these environments, but no clear trend emerges. For highway settings with bad weather, we can observe poor performance. As the pixel-wise anomaly scores are only calculated for bounding boxes and all other pixels are set to 0, there is generally a low false positive rate. However, if more than 5% of the ground truth anomalies are not detected, detecting all anomalous pixels requires a threshold close to zero. This results in the entire image being detected as anomalous, causing the FPR95 to rise sharply, as shown in Figure 5.
Table 3: Empirical study to determine the effects of different weights for HF2-VADAD.
$w_{r}$ and $w_{r_{p}}$ $w_{p}$ and $w_{p_{p}}$ FPR95 $\downarrow$ AUROC $\uparrow$
10 0.1 2.83 79.09
1 0.1 2.86 78.78
1 0 2.83 79.13
0 1 96.67 69.15
0.1 1 4.05 76.20
0.1 10 95.78 70.70
Figure 5: Comparison of ROC curves for the setting highway with rain. The x marks the FPR95 position. On the left, a low FPR95 is achieved with predicted bounding boxes. On the right, the FPR95 rose sharply with ground truth bounding boxes.
To determine optimal values for the weights, as introduced in Section 3, we performed an empirical study, as shown in Table 3. While we observe the best results in the case when we completely neglect image predictions, we decided to set $w_{r}$, $w_{r_{p}}=10$ and $w_{p}$, $w_{p_{p}}=0.1$, as the performance metrics were too small to be significant, and we wanted to retain the advantages of both modules. The study showed that the model performs significantly worse when image prediction is weighted higher, which is in contrast to the original implementation of HF2-VAD, in which the frame prediction was weighted higher than the flow reconstruction, with the weight for frame prediction set to 1 and the weight for flow reconstruction set to 0.1.
Ablation Studies. To better understand the performance of our approach, we perform three types of ablation studies. First, we only utilize the intermediate outputs of the ML-ML-MemAE-SC to evaluate isolated flow reconstructions.
Second, we evaluate both bounding boxes generated from an off-the-shelf object detection model [60] and ground truth bounding boxes from the evaluation dataset to isolate the effect of model failures. We can observe that bounding boxes are best detected in simpler highway settings with good weather. The IoU performance suffers especially from poor detections of distant objects. Since mostly irrelevant vehicle are missing in the detections, we see improved FPR95 values for bounding box detections, but at the cost of reduced AUROC scores.
Lastly, we examine the influence of the fine-tuning stage during training. After fine-tuning, HF2-VADAD shows an improvement in precision. While the AUROC decreases slightly, the model achieves a better balance in the performance metrics with a significantly lower FPR95.
4.2 Qualitative Results
Figure 6: On the left, two normal scenarios are shown. Here, the model’s next-frame predictions (a) match the actual next frames (b) closely. On the right, we have shown sudden brake scenarios. Here, significant prediction differences arise, leading to high pixel-wise anomaly values (c).
In Figure 6, we show a series of comparative visualizations demonstrating the performance of HF2-VADAD in traffic scenarios. The model is able to predict regular behaviors very well, while atypical ones lead to poor predictions and, thus, high anomaly scores. The pixel-wise anomaly maps serve as a meaningful indicator, with more highlighted areas signaling larger prediction discrepancies. For normal driving patterns, these maps show only small deviations from ground truth, indicating that the model can predict the subsequent images well. For anomalies, the model often predicts, as it learned, that the vehicle maintains the distance. This leads to positive anomaly detections, as the vehicle is currently anomalous and getting closer due to the braking maneuver. This can be recognized by the increased highlighting in the difference maps.
Failure Cases. As shown in Figure 7, there are certain cases which are challenging. The first image shows how the model often incorrectly produces high anomaly scores when a lead vehicle is turning rather than going straight. The middle image shows a scene where the object detection model has problems with detecting objects in unfavorable weather conditions, such as heavy rain so anomaly detection is not possible.
Figure 7: Failure Cases: The model can produce high pixel-wise anomaly scores for turning vehicles, as shown on the left. Here, the encircled area in the frame-wise graph shows the duration of the curve. In the middle, a missing bounding box detection during bad weather situations leads to no anomaly detection at all. On the right, the bounding box only captures a part of the lead vehicle, and a truck in the back is completely missed.
5 Conclusion
In summary, our research has shown that HF2-VAD, a framework originally developed for detecting anomalies in surveillance systems can be effectively transferred to the field of autonomous driving. By extending it do dense pixel-wise anomaly maps and creating a dedicated training and test dataset, we have validated the effectiveness of HF2-VADAD in the vehicle context, bridging the gap between static monitoring and dynamic vehicle environments.
Our contributions further include the extension to dense anomaly detections, domain-specific weight adaptions, and dealing with multiple bounding boxes separately rather than combining them into one score, using robust scaling. In the future, we aim to further refine the framework to better deal with the known failure cases and evaluate more types of anomalies to examine its performance among a wider variety of scenarios, such as wrong-way drivers.